OpenStudy (anonymous):

What is the slope of a line containing the points (−2, −4) and (4, 5)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

(5 - (-4))/(4 - (-2)) = 9/6 = 3/2